# ADAPT to Robustify Prompt Tuning Vision Transformers

### v1

@@ 1 Introduction
Despite their success, deep neural networks, including the popular ViT architecture, have been shown to be vulnerable to adversarial attacks [27, 24]. These attacks are modifications to input images, which are generally imperceptible to humans but can drastically change the prediction of a neural network. Many countermeasures have been introduced to induce robustness against such attacks, one common approach being adversarial training [23], where the input is augmented with said perturbations during training. While originally introduced for CNNs, adversarial training has been extended to ViTs in the full-model fine-tuning paradigm [25]. However, adversarial training and similar methods are computationally expensive and difficult to perform for large datasets as it requires performing multi-step evaluation of adversarial examples. Hence ViTs are generally pre-trained without such considerations.
Transformer-based language models [2] which can consist of billions of parameters require pretraining on massive amounts of data. To fine-tune such models for different tasks, one would have to save a separate copy of the weights per task, making it highly inefficient in terms of parameters. However, it is shown that these models can be adapted to downstream tasks with fewer parameters, for example using prompt design [2] or prompt tuning [19]. In prompt tuning, optimizable tokens are prepended to the model input in order to leverage the pretraining knowledge and adapt to the downstream task without changing the weights of the model. This allows for a lightweight adaptation of these large models. ViTs can benefit from the same technique[15], especially as their parameters continue to increase [8].
It is therefore important to explore the utility of prompt tuning for the adversarial robustness of ViTs. In this paper, we show that existing adversarial defense methods, when applied to the prompt tuning paradigm, suffer from gradient obfuscation [1]. This means the gradients of the model are not useful for obtaining adversarial perturbations, and training using existing methods will lead to a false sense of security. We demonstrate this phenomenon empirically using single-step attacks, and then design an adaptive attack to expose the vulnerability of existing methods. Finally, we propose a framework for ADaptive Adversarial Prompt Tuning (ADAPT) to overcome this vulnerability. We empirically demonstrate the superior robustness of our method in multiple settings.
@@
  • We investigate the adversarial robustness of the prompt tuning paradigm for ViTs. We show that the existing methods, when applied to the prompt tuning scenario, suffer from gradient obfuscation [1].
@@
(b) ADAPT-CE + Prompt Tuning
Figure 1: Existing methods do not exhibit robustness to adaptive attacks. Comparison of the cross entropy loss of a random batch along two adversarial perturbation directions. Left (a) depicts the loss values of a prompt trained with traditional adversarial training and right (b) shows the loss values of a prompt trained with ADAPT-CE. Adversarial Training + Prompt Tuning does not exhibit a significant increase in the loss under the traditional PGD direction, but shows significant vulnerability to adaptive attacks. In contrast, the proposed method ADAPT-CE, exhibits robustness to both perturbation directions.
@@ 2 Related Works
Vision Transformers. Inspired by the groundbreaking success of Transformers for Natural Language Processing (NLP) tasks [29, 16, 26], Vision Transformers (ViTs) [9] are an adaptation of the architecture for computer vision tasks. Similar to the ‘tokens’ fed to NLP transformers, here the input images are split into fixed-size patches, which are then embedded and fed to the transformer layers of the ViT. ViTs are powerful architectures and achieve state-of-the-art performance on several downstream vision tasks [28, 5, 12].
Prompting. This technique was introduced in the NLP literature to eliminate the need to fine-tune large pre-trained language models for downstream tasks [21]. Intuitively, prompting reformulates the downstream task so that it closely resembles the task that a frozen model has been pre-trained to solve. Prompt tuning [19], i.e., training the prompts for downstream tasks, has been extended to ViTs and is highly effective for domain generalization [35], continual learning [30], and several other vision tasks.
Adversarial Robustness. Adversarial Training (AT) was first introduced for CNNs and was thoroughly investigated [27, 11, 23] with several variations [3, 33, 34]. ViTs have also been shown to be vulnerable to adversarial attacks [24]. However, AT and its variants are computationally expensive due to requiring multiple forward and backward passes to obtain adversarial examples to train with. This makes them difficult to use for pre-training ViTs. Therefore, adversarial defense methods are typically only adopted during fine-tuning for downstream tasks [25].
Robust Prompts. Prompt tuning has been employed for enhancing the robustness of NLP transformers [32], which deal with discrete inputs. Extending the same robustness to the continuous input space for ViTs is non-trivial and has not been explored in the literature.
There are other works which explore robustness of visual prompt tuning [6]. However, they are not tailored to the transformer paradigm. Moreover, their number of tuned parameters grows with the number of classes in the dataset. Visual prompting or adversarial reprogramming was proposed before the advent of prompt tuning for adapting CNNs to downstream tasks [10]. As such, we are the first to investigate robust prompting for vision transformers.
@@ 3.1 Vision Transformer
The Vision Transformer [9] has proven to be quite effective at performing various computer vision tasks. For our purposes, we consider the classification task. A transformer $f_{\Phi,\Psi}$ as a classifier consists of a feature extractor $\Phi$ and a linear classifier $\Psi$ operating on the extracted features. That is,
@@ 3.2 Prompt Tuning
With the advent of transformers and large language models with billions of parameters pre-trained on extremely large amounts of data, it is desirable to tune these models for downstream tasks at a minimal cost. While full-model fine-tuning is an effective option, it requires training and storing an entirely new set of model weights for each downstream task, which becomes inefficient in terms of memory as both the number of tasks and the number of model parameters grow. A simple yet effective solution to this is prompt tuning (PT) [19], which proceeds by simply appending tunable prompt tokens to the transformer input space. Depending on the model size and number of tokens, it can yield comparable performance to full-model fine-tuning. This allows us to adapt to downstream tasks without having to store copies of extremely large models, and instead only storing parameter-efficient prompts.
@@
where [.,.][.,.][ . , . ] is the concatenation operator, $d$ is the hidden dimension of the transformer, and $m$ is the prompt length (recall that the input space of the transformer is arbitrary length sequences). Another way to perform prompt tuning is Prefix tuning [20] or PT-v2 tuning (PT2) [22], where similar to PT, the backbone weights are kept unchanged. However, a set of prompt tokens are instead directly prepended to the input of every layer of the transformer, by only prepending to the key and value matrices of the self-attention mechanism (while keeping the same number of tokens per layer). That is, $f(\theta_{p},x)$ in the case of PT2 is as follows
@@ 3.3 Adversarial Robustness
The adversarial example defined above is an optimization problem that is typically solved using methods such as Projected Gradient Descent (PGD) [23]. As deep learning models are highly susceptible to such norm-bounded perturbations [27], it is crucial to defend against these attacks. However, it is important to note that PGD assumes full knowledge of the model parameters and inference algorithm, i.e. a white-box setting, for the adversary. Adversarial robustness can also be investigated through the lens of adversaries with limited knowledge, namely the black-box [14], and grey-box [31] settings. In this work, we focus on the white-box scenario.
One widely used defense method against adversarial attacks is Adversarial Training (AT) [23]. If $f_{\theta}$ is a classifier parameterized by $\theta$, AT tries to solve the following minimax optimization problem:
@@
An adversarial defense method suffers from gradient masking or obfuscated gradients [1] when it does not have "useful gradients" for generating adversarial examples. Typically, there are three main causes for gradient obfuscation: exploding and vanishing gradients, stochastic gradients, and shattered gradients. We show that traditional adversarial defense methods in the prompt tuning paradigm suffer from shattered gradients. This means that the gradient of the model w.r.t. the input is incorrect and does not properly maximize the classification loss. To validate this claim, we use the observation of [1] that in the presence of obfuscated gradients, single-step attacks are typically more effective than their multi-step counterparts.
@@ 4.1 Adaptive Adversarial Attack
In the prompt tuning scenario, $\nabla_{x}f(x)$ does not necessarily equal $\nabla_{x}f(\theta_{p},x)$ Therefore, the existence of the prompt may lead to a form of gradient obfuscation [1], giving us a false sense of security (See Fig. 0(a)). In the experiment section (Sec. 5.2), we demonstrate this effect by designing an adaptive attack and empirically show that while previous adversarial training methods may be robust against traditional attacks, they are vulnerable to our adaptive attack which reduces their robust accuracy to nearly zero.
@@
We can solve this optimization problem using Projected Gradient Descent [23] similar to Eq. 11. Here, the key difference lies in conditioning on the prompt during each gradient step.
@@ 4.2 Prompt Adversarial Training
Cross Entropy (CE). We can choose the cross-entropy loss for the prediction of the adaptive adversarial example:
@@
KL Divergence. We can choose $\mathcal{L}_{adv}$ to be the KL-divergence between the prediction distributions of the unperturbed and perturbed samples:
@@ 4.3 Additional Design Choices
Prompting the feature extractor. Prompt tuning can be performed in a multitude of ways by changing the definition of $f(\theta_{p},x)$. In our experiments, we observe that for adversarial robustness, using PT2 (otherwise known as Prefix Tuning) significantly outperforms using only PT. This behavior is consistently observed when using the same training scheme and total number of tuned parameters across the two methods. We hypothesize that this performance gap exists since the feature extractor $\Phi^{*}$ of the frozen backbone is not adversarially robust. Since PT simply prepends additional tokens only to the input of the frozen feature extractor, this results in the same function $\Phi^{*}$ on a different set of inputs. However, PT2 prepends prompt tokens directly in each layer of the feature extractor, resulting in a modified feature extractor $\hat{\Phi}$.
Tuning the embedding layer. Traditionally, the prompt tokens are introduced after the embedding function $T$ and do not influence the mapping of the image patches to the embedding space. That is, the patch embedding remains unchanged by prompt tuning, which can cause vulnerability. Since the patch embedding does not have a significantly large number of parameters compared to the frozen backbone, we opt to tune the patch embedding parameters $T$ along with the prompt tokens $\theta_{p}$ and the linear classifier head $\Psi$.
Algorithm 1 shows a pseudocode of our training algorithm. We provide our code as supplementary material, and we will release a public version of our code upon acceptance.
@@ 5.1 Experimental Setting
Datasets. We perform experiments on CIFAR-10[17], CIFAR100 [18], and Imagenette [13]. Note that these datasets are widely employed benchmarks for evaluating adversarial defenses since training and evaluation on ImageNet using multi-step gradient-based attacks requires significant computational time and resources.
Adversarial Robustness Methods. We compare ADAPT to the following state-of-the-art methods established in the literature:
  • Adversarial Training [23]: As described in Eq. 12, Adversarial Training solves a minimax optimization objective by augmenting the inputs during training using traditional adversarial attacks.
  • TRADES [33]: TRADES identifies a trade-off between robustness and accuracy. It designs a defense method accordingly, which performs training with a modified loss function, performing natural training and at the same time reducing the prediction distance of unperturbed samples with that of samples perturbed to maximize the same distance.
Architectures and Tuned Parameters. We perform most of our experiments on the ViT-B model. For prompt tuning (PT), we keep the number of tuned parameters within $\sim 1\%$ of the total model parameters. We provide additional results on the ViT-S and ViT-L model configurations.
As discussed in Sec. 4.3, due to superior performance, we perform most experiments with PT2 and tune the patch embedding as well as the linear classifier (all of which we take into account into the number of tuned parameters). We also perform ablation studies on PT while freezing the patch embedding layer.
Evaluation Metrics. We evaluate each method with regards to accuracy on the test set. We evaluate on unperturbed samples, also known as clean accuracy, as well as perturbed samples. In Sec. 5.2, we use traditional adversarial attacks as well as adaptive ones to validate our claims of gradient obfuscation. In all other sections, our attacks are adaptive to ensure a fair worst-case evaluation. We use adaptive PGD10 Eq. 13, adaptive CW [4], which is obtained by replacing the loss in Eq. 13 with the CW loss, as well as AutoAttack (AA) [7], an ensemble of diverse and parameter free white-box and black-box attacks.

### v2

@@ 1 Introduction
Despite their success, deep neural networks, including the popular ViT architecture, have been shown to be vulnerable to adversarial attacks (Szegedy et al., 2014; Mahmood et al., 2021). These attacks are modifications to input images, which are generally imperceptible to humans but can drastically change the prediction of a neural network. Many countermeasures have been introduced to induce robustness against such attacks, one common approach being adversarial training (Madry et al., 2018), where the input is augmented with said perturbations during training. While originally introduced for CNNs, adversarial training has been extended to ViTs in the full-model fine-tuning paradigm (Mo et al., 2022). However, adversarial training and similar methods are computationally expensive and difficult to perform for large datasets as it requires performing multi-step evaluation of adversarial examples. Hence ViTs are generally pre-trained without such considerations.
Transformer-based language models (Brown et al., 2020) which can consist of billions of parameters require pretraining on massive amounts of data. To fine-tune such models for different tasks, one would have to save a separate copy of the weights per task, making it highly inefficient in terms of parameters. However, it is shown that these models can be adapted to downstream tasks with fewer parameters, for example using prompt design (Brown et al., 2020) or prompt tuning (PT) (Lester et al., 2021). In prompt tuning, optimizable tokens are prepended to the model input in order to leverage the pretraining knowledge and adapt to the downstream task without changing the weights of the model. This allows for a lightweight adaptation of these large models. ViTs can benefit from the same technique (Jia et al., 2022), especially as their parameters continue to increase (Dehghani et al., 2023).
It is therefore important to explore the utility of prompt tuning for the adversarial robustness of ViTs. In this paper, we show that existing adversarial defense methods, when applied to the prompt tuning paradigm, suffer from gradient obfuscation (Athalye et al., 2018). This means the gradients of the model are not useful for obtaining adversarial perturbations, and training using existing methods will lead to a false sense of security. We demonstrate this phenomenon empirically using single-step attacks, and then design an adaptive attack to expose the vulnerability of existing methods. Finally, we propose a framework for ADaptive Adversarial Prompt Tuning (ADAPT) to overcome this vulnerability. We empirically demonstrate the superior robustness of our method in multiple settings.
@@
  • We investigate the adversarial robustness of the prompt tuning paradigm for ViTs. We show that the existing methods, when applied to the prompt tuning scenario, suffer from gradient obfuscation (Athalye et al., 2018).
@@
Our implementation is available on Github [^1] https://github.com/Gnomy17/robust_prompt.
@@
(b) $\text{ADAPT}_{CE}$ + Prompt Tuning
Figure 1: Existing methods do not exhibit robustness to adaptive attacks. Comparison of the cross entropy loss of a random batch along two adversarial perturbation directions. Left (a) depicts the loss values of a prompt trained with traditional adversarial training and right (b) shows the loss values of a prompt trained with $\text{ADAPT}_{CE}$. Adversarial Training + Prompt Tuning does not exhibit a significant increase in the loss under the traditional PGD direction, but shows significant vulnerability to adaptive attacks. In contrast, the proposed method $\text{ADAPT}_{CE}$, exhibits robustness to both perturbation directions.
@@ 2 Related Works
Vision Transformers.
Inspired by the groundbreaking success of Transformers for Natural Language Processing (NLP) tasks (Vaswani et al., 2017; Kenton & Toutanova, 2019; OpenAI, 2023), Vision Transformers (ViTs) (Dosovitskiy et al., 2020) are an adaptation of the architecture for computer vision tasks. Similar to the ‘tokens’ fed to NLP transformers, here the input images are split into fixed-size patches, which are then embedded and fed to the transformer layers of the ViT. ViTs are powerful architectures and achieve state-of-the-art performance on several downstream vision tasks (Touvron et al., 2021; Caron et al., 2021; He et al., 2022).
Prompting.
This technique was introduced in the NLP literature to eliminate the need to fine-tune large pre-trained language models for downstream tasks (Liu et al., 2023). Initially, Prompting was used to fine-tune the behavior of language models by pre-pending a set of fixed prompt templates to the input text. Intuitively, prompting reformulates the downstream task so that it closely resembles the task that a frozen model has been pre-trained to solve. For example, Gao et al. (2021) use prompting for few-shot learning of language models. Related to this work, Raman et al. (2023) investigate the effects of adversarial robustness of prompting language models. However, these works are fundamentally specific to the NLP literature, as they require pre-determined prompts, and investigate the problem in the text input domain which is discrete. The extension of fixed prompt templates to the vision space, where inputs are continuous is non-trivial. To that effect, prompt tuning was introduced (Lester et al., 2021) which uses continuous tunable vectors in the embedding space as prompts. Prompt-tuning, i.e., training the prompts for downstream tasks, has been extended to ViTs and is highly effective for domain generalization (Zheng et al., 2022), continual learning (Wang et al., 2022), and several other vision tasks.
Adversarial Robustness.
Adversarial Training (AT) was first introduced for CNNs and was thoroughly investigated (Szegedy et al., 2014; Goodfellow et al., 2015; Madry et al., 2018) with several variations (Buckman et al., 2018; Zhang et al., 2019; 2020). ViTs have also been shown to be vulnerable to adversarial attacks (Mahmood et al., 2021). However, AT and its variants are computationally expensive due to requiring multiple forward and backward passes to obtain adversarial examples to train with. This makes them difficult to use for pre-training ViTs. Therefore, adversarial defense methods are typically only adopted during fine-tuning for downstream tasks (Mo et al., 2022).
Robust Prompts.
Prompt tuning has been employed for enhancing the robustness of NLP transformers (Yang et al., 2022), which deal with discrete inputs. Extending the same robustness to the continuous input space for ViTs is non-trivial and has not been explored in the literature.
There are other works which explore robustness of visual prompt tuning (Chen et al., 2023a). However, they are not tailored to the transformer paradigm. Moreover, their number of tuned parameters grows with the number of classes in the dataset. Visual prompting or adversarial reprogramming was proposed before the advent of prompt tuning for adapting CNNs to downstream tasks (Elsayed et al., 2019). As such, we are the first to investigate robust prompting for vision transformers.
@@ 3.1 Vision Transformer
The Vision Transformer (Dosovitskiy et al., 2020) has proven to be quite effective at performing various computer vision tasks. For our purposes, we consider the classification task. A transformer $f_{\Phi,\Psi}$ as a classifier consists of a feature extractor $\Phi$ and a linear classifier $\Psi$ operating on the extracted features. That is,
@@ 3.2 Prompt Tuning
With the advent of transformers and large language models with billions of parameters pre-trained on extremely large amounts of data, it is desirable to tune these models for downstream tasks at a minimal cost. While full-model fine-tuning is an effective option, it requires training and storing an entirely new set of model weights for each downstream task, which becomes inefficient in terms of memory as both the number of tasks and the number of model parameters grow. A simple yet effective solution to this is prompt tuning (Lester et al., 2021), which proceeds by simply appending tunable prompt tokens to the transformer input space. Depending on the model size and number of tokens, it can yield comparable performance to full-model fine-tuning. This allows us to adapt to downstream tasks without having to store copies of extremely large models, and instead only storing parameter-efficient prompts.
@@
where [.,.][.,.][ . , . ] is the concatenation operator, $d$ is the hidden dimension of the transformer, and $m$ is the prompt length (recall that the input space of the transformer is arbitrary length sequences). Another way to perform prompt tuning is Prefix tuning (Li & Liang, 2021) or PT-v2 tuning (PT2) (Liu et al., 2022), where similar to PT, the backbone weights are kept unchanged. However, a set of prompt tokens are instead directly prepended to the input of every layer of the transformer, by only prepending to the key and value matrices of the self-attention mechanism (while keeping the same number of tokens per layer). That is, $f(\theta_{p},x)$ in the case of PT2 is as follows
@@ 3.3 Adversarial Robustness
The adversarial example defined above is an optimization problem that is typically solved using methods such as Projected Gradient Descent (PGD) (Madry et al., 2018). As deep learning models are highly susceptible to such norm-bounded perturbations (Szegedy et al., 2014), it is crucial to defend against these attacks. However, it is important to note that PGD assumes full knowledge of the model parameters and inference algorithm, i.e. a white-box setting, for the adversary. Adversarial robustness can also be investigated through the lens of adversaries with limited knowledge, namely the black-box (Ilyas et al., 2018), and grey-box (Xu et al., 2021) settings. In this work, we focus on the white-box scenario.
One widely used defense method against adversarial attacks is Adversarial Training (Madry et al., 2018). If $f_{\theta}$ is a classifier parameterized by $\theta$, AT tries to solve the following minimax optimization problem:
@@
An adversarial defense method suffers from gradient masking or obfuscated gradients (Athalye et al., 2018) when it does not have "useful gradients" for generating adversarial examples. Typically, there are three main causes for gradient obfuscation: exploding and vanishing gradients, stochastic gradients, and shattered gradients. We show that traditional adversarial defense methods in the prompt tuning paradigm suffer from shattered gradients. This means that the gradient of the model w.r.t. the input is incorrect and does not properly maximize the classification loss. To validate this claim, we use the observation of (Athalye et al., 2018) that in the presence of obfuscated gradients, single-step attacks are typically more effective than their multi-step counterparts.
@@ 4.1 Adaptive Adversarial Attack
In the prompt tuning scenario, $\nabla_{x}f(x)$ does not necessarily equal $\nabla_{x}f(\theta_{p},x)$ Therefore, the existence of the prompt may lead to a form of gradient obfuscation (Athalye et al., 2018), giving us a false sense of security (See fig. 1(a)). In the experiment section , we demonstrate this effect by designing an adaptive attack and empirically show that while previous adversarial training methods may be robust against traditional attacks, they are vulnerable to our adaptive attack which reduces their robust accuracy to nearly zero.
@@
We can solve this optimization problem using Projected Gradient Descent (Madry et al., 2018) similar to eq. 11. Here, the key difference lies in conditioning on the prompt during each gradient step.
@@ 4.2 Prompt Adversarial Training
Cross Entropy (CE).
We can choose the cross-entropy loss for the prediction of the adaptive adversarial example:
@@
KL Divergence.
We can choose $\mathcal{L}_{adv}$ to be the KL-divergence between the prediction distributions of the unperturbed and perturbed samples:
@@ 4.3 Additional Design Choices
Prompting the feature extractor.
Prompt tuning can be performed in a multitude of ways by changing the definition of $f(\theta_{p},x)$. In our experiments, we observe that for adversarial robustness, using PT2 (otherwise known as Prefix Tuning) significantly outperforms using only PT. This behavior is consistently observed when using the same training scheme and total number of tuned parameters across the two methods. We hypothesize that this performance gap exists since the feature extractor $\Phi^{*}$ of the frozen backbone is not adversarially robust. Since PT simply prepends additional tokens only to the input of the frozen feature extractor, this results in the same function $\Phi^{*}$ on a different set of inputs. However, PT2 prepends prompt tokens directly in each layer of the feature extractor, resulting in a modified feature extractor $\hat{\Phi}$.
Tuning the embedding layer.
Traditionally, the prompt tokens are introduced after the embedding function $T$ and do not influence the mapping of the image patches to the embedding space. That is, the patch embedding remains unchanged by prompt tuning, which can cause vulnerability. Since the patch embedding does not have a significantly large number of parameters compared to the frozen backbone, we opt to tune the patch embedding parameters $T$ along with the prompt tokens $\theta_{p}$ and the linear classifier head $\Psi$.
Algorithm 1 shows a pseudocode of our training algorithm. We provide our code as supplementary material, and we will release an open-source version of our code upon acceptance.
@@ 5.1 Experimental Setting
Datasets.
We perform experiments on CIFAR-10(Krizhevsky et al., a), CIFAR100 (Krizhevsky et al., b), and Imagenette (Howard, ). Note that these datasets are widely employed benchmarks for evaluating adversarial defenses since training and evaluation on ImageNet using multi-step gradient-based attacks requires significant computational time and resources.
Adversarial Robustness Methods.
We compare ADAPT to the following state-of-the-art methods established in the literature:
  • Adversarial Training (Madry et al., 2018): As described in eq. 12, Adversarial Training solves a minimax optimization objective by augmenting the inputs during training using traditional adversarial attacks.
  • TRADES (Zhang et al., 2019): TRADES identifies a trade-off between robustness and accuracy. It designs a defense method accordingly, which performs training with a modified loss function, performing natural training and at the same time reducing the prediction distance of unperturbed samples with that of samples perturbed to maximize the same distance.
  • MART (Wang et al., 2019): MART builds on top of TRADES and emphasizes on misclassified examples during training to improve robustness.
  • NFGSM (de Jorge Aranda et al., 2022) NFGSM proposes to use a high-magnitude noise perturbation as an initialization point for generating adversarial attacks during single-step adversarial training to avoid overfitting
Architectures and Tuned Parameters.
We perform most of our experiments on the ViT-B model. For prompt tuning (PT), we keep the number of tuned parameters within $\sim 1\%$ of the total model parameters. We provide additional results on the ViT-S and ViT-L model configurations.
As discussed in design choices, due to superior performance, we perform most experiments with PT2 and tune the patch embedding as well as the linear classifier (all of which we take into account into the number of tuned parameters). We also perform ablation studies on PT while freezing the patch embedding layer.
Evaluation Metrics.
We evaluate each method with regards to accuracy on the test set. We evaluate on unperturbed samples, also known as clean accuracy, as well as perturbed samples. In one experiment, as mentioned, we use traditional adversarial attacks as well as adaptive ones to validate our claims of gradient obfuscation. In all other sections, excluding the black-box attacks section, our attacks are white-box and adaptive to ensure a fair worst-case evaluation. We use adaptive PGD10 eq. 13, adaptive CW (Carlini & Wagner, 2017), which is obtained by replacing the loss in eq. 13 with the CW loss, as well as AutoAttack (AA) (Croce & Hein, 2020), an ensemble of diverse and parameter free white-box and black-box attacks. Finally, we evaluate against two state of the art blackbox attacks: RayS (Chen & Gu, 2020) and AdaEA (Chen et al., 2023b).